\vspace*{0cm}





\begin{thebibliography}{99}

\bibitem{KamiyaSato1}N. Kamiya, M. Sato, ``Hermitian generalised Jordan triple systems and certain applications to field theory," preprint.

\bibitem{KamiyaSato2}N.Kamiya and M.Sato, ``Hermitian $(\epsilon, \delta)$-Freudenthal-Kantor triple systems and certain applications of *-generalized Jordan triple systems to field theory," to appear in AHEP 


\bibitem{modules1}S. Palmer, C. Saemann, ``M-brane Models from Non-Abelian Gerbes," JHEP 07 (2012) 010.

\bibitem{modules2}S. Palmer, C. Saemann, ``Six-Dimensional (1,0) Superconformal Models and Higher Gauge Theory," J. Math. Phys. 54 (2013) 113509.

\bibitem{modules3}S. Palmer, C. Saemann, ``The ABJM Model is a Higher Gauge Theory," arXiv:1311.1997.

%\bibitem{}, ``," {\bf }.








\bibitem{Nambu}Y. Nambu, Generalized Hamiltonian dynamics, Phys. Rev.  D7 (1973) 2405.


\bibitem{Filippov}V. T. Filippov, n-Lie algebras, Sib. Mat. Zh. 26, No. 6, (1985) 126140.



\bibitem{BergshoeffSezginTownsend}E. Bergshoeff, E. Sezgin, P.K. Townsend, Supermembranes and Eleven-Dimensional Supergravity, Phys. Lett.  B189 (1987) 75.


\bibitem{deWHN}B. de Wit, J. Hoppe, H. Nicolai, On the Quantum Mechanics of Supermembranes, Nucl. Phys.  B305 (1988) 545.



\bibitem{Yoneya}H. Awata, M. Li, D. Minic, T. Yoneya, On the Quantization of Nambu Brackets, JHEP  0102 (2001) 013.

\bibitem{Minic}D. Minic, ``M-theory and Deformation Quantization," arXiv:hep-th/9909022.








\bibitem{BLG1}J. Bagger, N. Lambert, ``Modeling Multiple M2's," Phys. Rev.  D75: 045020, 2007, hep-th/0611108.

\bibitem{BLG2}A. Gustavsson, ``Algebraic structures on parallel M2-branes," Nucl. Phys.  B811: 66, 2009, arXiv:0709.1260 [hep-th].

\bibitem{BLG3}J. Bagger, N. Lambert, ``Gauge Symmetry and Supersymmetry of Multiple M2-Branes," Phys. Rev. D77: 065008, 2008, arXiv:0711.0955 [hep-th].

\bibitem{BLG4}O. Aharony, O. Bergman, D. L. Jafferis, J. Maldacena, ``N=6 superconformal Chern-Simons-matter theories, M2-branes and their gravity duals," JHEP 0810: 091, 2008, arXiv:0806.1218 [hep-th]

\bibitem{BLG5}J. Bagger, N. Lambert, ``Three-Algebras and N=6 Chern-Simons Gauge Theories," Phys. Rev.  D79: 025002, 2009, arXiv:0807.0163 [hep-th]





\bibitem{Nogo1}J. Figueroa-O'Farrill, G. Papadopoulos, ``Pluecker-type relations for orthogonal planes," J. Geom. Phys. 49 (2004) 294, math/0211170]

\bibitem{Nogo2}G. Papadopoulos, ``M2-branes, 3-Lie Algebras and Plucker relations," JHEP 0805 (2008) 054, arXiv:0804.2662 [hep-th]

\bibitem{Nogo3}J. P. Gauntlett, J. B. Gutowski, ``Constraining Maximally Supersymmetric Membrane Actions," JHEP 0806 (2008) 053, arXiv:0804.3078 [hep-th] 





\bibitem{sp1}D. Gaiotto, E. Witten, ``Janus Configurations, Chern-Simons Couplings, And The Theta-Angle in N=4 Super Yang-Mills Theory," arXiv:0804.2907[hep-th]

\bibitem{Iso}Y. Honma, S. Iso, Y. Sumitomo, S. Zhang, ``Janus field theories from multiple M2 branes," Phys.Rev. D78 :025027,2008, arXiv:0805.1895 [hep-th].

\bibitem{sp2}K. Hosomichi, K-M. Lee, S. Lee, S. Lee, J. Park, ``N=5,6 Superconformal Chern-Simons Theories and M2-branes on Orbifolds," JHEP 0809 (2008) 002, arXiv:0806.4977[hep-th]

\bibitem{sp3}M. Schnabl, Y. Tachikawa, ``Classification of N=6 superconformal theories of ABJM type," arXiv:0807.1102[hep-th]




\bibitem{Lorentz0}S. Mukhi, C. Papageorgakis, M2 to D2, JHEP  0805 (2008) 085.


\bibitem{Nogo4}J. Gomis, G. Milanesi, J. G. Russo, ``Bagger-Lambert Theory for General Lie Algebras," JHEP  0806: 075, 2008, arXiv:0805.1012 [hep-th].

\bibitem{Nogo5}S. Benvenuti, D. Rodriguez-Gomez, E. Tonni, H. Verlinde, ``N=8 superconformal gauge theories and M2 branes," JHEP 0901: 078, 2009, arXiv:0805.1087 [hep-th].

\bibitem{Nogo6}P.-M. Ho, Y. Imamura, Y. Matsuo, ``M2 to D2 revisited," JHEP 0807: 003, 2008, arXiv:0805.1202 [hep-th].

\bibitem{Nogo7}M. A. Bandres, A. E. Lipstein, J. H. Schwarz, ``Ghost-Free Superconformal Action for Multiple M2-Branes," JHEP 0807: 117, 2008, arXiv:0806.0054 [hep-th]

\bibitem{text}P. de Medeiros, J. Figueroa-O'Farrill, E. Me'ndez-Escobar, P. Ritter, ``On the Lie-algebraic origin of metric 3-algebras," Commun. Math. Phys. 290 (2009) 871, arXiv:0809.1086 [hep-th]

\bibitem{Nogo8}S. A. Cherkis, V. Dotsenko, C. Saeman, ``On Superspace Actions for Multiple M2-Branes, Metric 3-Algebras and their Classification," Phys. Rev. D79 (2009) 086002, arXiv:0812.3127 [hep-th]


\bibitem{Nogo9}P.-M. Ho, Y. Matsuo, S. Shiba, ``Lorentzian Lie (3-)algebra and toroidal compactification of M/string theory," arXiv:0901.2003 [hep-th]

\bibitem{Nogo10}P. de Medeiros, J. Figueroa-O'Farrill, E. Mendez-Escobar, P. Ritter, ``Metric 3-Lie algebras for unitary Bagger-Lambert theories," JHEP 0904: 037, 2009, arXiv:0902.4674 [hep-th]







\bibitem{NishinoRajpoot}H. Nishino, S. Rajpoot, Triality and Bagger-Lambert Theory, Phys. Lett.  B671 (2009) 415,  arXiv:0901.1173.

\bibitem{GustavssonRey}A. Gustavsson, S-J. Rey, Enhanced N=8 Supersymmetry of ABJM Theory on R(8) and R(8)/Z(2), arXiv:0906.3568 [hep-th].

\bibitem{ABJ}O. Aharony, O. Bergman, D. L. Jafferis, Fractional M2-branes, JHEP  0811 (2008) 043,  arXiv:0807.4924.

\bibitem{HanadaMannelliMatsuo}M. Hanada, L. Mannelli, Y. Matsuo, Large-N reduced models of supersymmetric quiver, Chern-Simons gauge theories and ABJM, arXiv:0907.4937 [hep-th].

\bibitem{IshikiShimasakiTsuchiya}G. Ishiki, S. Shimasaki, A. Tsuchiya, Large N reduction for Chern-Simons theory on $S^3$, Phys. Rev.  D80 (2009) 086004,  arXiv:0908.1711.

\bibitem{KawaiShimasakiTsuchiya}H. Kawai, S. Shimasaki, A. Tsuchiya, Large N reduction on group manifolds, arXiv:0912.1456 [hep-th].

\bibitem{IshikiShimasakiTsuchiya2}G. Ishiki, S. Shimasaki, A. Tsuchiya, A Novel Large-N Reduction on $S^3$: Demonstration in Chern-Simons Theory, arXiv:1001.4917 [hep-th].


\bibitem{PangWang}Y. Pang, T. Wang, From N M2's to N D2's, Phys. Rev.  D78 (2008) 125007,  arXiv:0807.1444.



\bibitem{DeBellisSaemannSzabo}J. DeBellis, C. Saemann, R. J. Szabo, Quantized Nambu-Poisson Manifolds in a 3-Lie Algebra Reduced Model, JHEP  1104 (2011) 075,  arXiv:1012.2236.



\bibitem{Jabbari}M. M. Sheikh-Jabbari, Tiny Graviton Matrix Theory: DLCQ of IIB Plane-Wave String Theory, A Conjecture , JHEP  0409 (2004) 017,  arXiv:hep-th/0406214.

\bibitem{GomisSalimPasserini}J. Gomis, A. J. Salim, F. Passerini, Matrix Theory of Type IIB Plane Wave from Membranes, JHEP  0808 (2008) 002,  arXiv:0804.2186.


\bibitem{HosomichiLeeLee}K. Hosomichi, K. Lee, S. Lee, Mass-Deformed Bagger-Lambert Theory and its BPS Objects, Phys.Rev.  D78 (2008) 066015,  arXiv:0804.2519.











\bibitem{Smolin1}L. Smolin, ``M theory as a matrix extension of Chern-Simons theory," Nucl.Phys. B591 (2000) 227, hep-th/0002009.
 
\bibitem{Smolin2}L. Smolin, ``The cubic matrix model and a duality between strings and loops," hep-th/0006137.

\bibitem{Azuma}T. Azuma, S. Iso, H. Kawai, Y. Ohwashi, ``Supermatrix Models," Nucl.Phys. B610 (2001) 251, hep-th/0102168.


\bibitem{3-graded}J. Palmkvist, ``Three-algebras, triple systems and 3-graded Lie superalgebras," J.Phys. A43 (2010) 015205, arXiv:0905.2468 [hep-th]

\bibitem{5-graded}S. Kim, J. Palmkvist, ``N=5 three-algebras and 5-graded Lie superalgebras," J.Math.Phys. 52 (2011) 083502, arXiv:1010.1457 [hep-th]

\bibitem{Palmkvist}J. Palmkvist, ``Unifying N = 5 and N = 6," JHEP 1105 (2011) 088, arXiv:1103.4860.





\bibitem{Nilsson}B.E.W. Nilsson, J. Palmkvist, ``Superconformal M2-branes and generalized Jordan triple systems," Class.Quant.Grav. 26 (2009) 075007, arXiv:0807.5134 [hep-th]



\bibitem{Chowdhury}S. P. Chowdhury, S. Mukhopadhyay, K. Ray, ``BLG theory with generalized Jordan triple systems," arXiv:0903.2951 [hep-th] 


\bibitem{bosonicM}M. Sato, `` Covariant Formulation of M-Theory," Int. J. Mod. Phys. A24 (2009), 5019, arXiv:0902.1333 [hep-th]

\bibitem{MModel}M. Sato, ``Model of M-theory with Eleven Matrices," JHEP 1007 (2010) 026, arXiv:1003.4694 [hep-th] 

\bibitem{LorentzianM}M. Sato, ``Supersymmetry and the Discrete Light-Cone Quantization Limit of the Lie 3-algebra Model of M-theory," Phys. Rev. D85 (2012) 046003, arXiv:1110.2969 [hep-th]

\bibitem{Zariski}M. Sato, ``Zariski Quantization as Second Quantization," Phys. Rev. D85 (2012) 126012, arXiv:1202.1466 [hep-th]


\bibitem{book}M. Sato, ``3-Algebras in String Theory," Linear Algebra - Theorems and Applications, Edited by Hassan Abid Yasser, InTech, Croatia, 2012

\bibitem{3BFSS}M. Sato, ``Three-Algebra BFSS Matrix Theory," Int.J.Mod.Phys. A28 (2013) 1350155, arXiv:1304.4430 [hep-th]
 	

\bibitem{3IIB}M. Sato, ``Extension of IIB Matrix Model by Three-Algebra," Int.J.Mod.Phys. A28 (2013) 1350083, arXiv:1304.4796 [hep-th]


\bibitem{4IIB}M. Sato, ``Four-algebraic extension of the IIB matrix model," PTEP 2013 (2013) 7, 073 B04, arXiv:1304.7904 [hep-th]
 


%\bibitem{}, ``," {\bf }, arXiv:





\bibitem{3} 
N.Kamiya,;A structure theory of 
Freudenthal-Kantor triple 
systems,J.Alg.
110
(1987)
108-123.
\bibitem{4} 
N.Kamiya,;A structure theory of Freudenthal-Kantor triple systems 
II,Comm.
Math.Univ.Sancti Pauli.,38(1989)41-60.
\bibitem{5}
N.Kamiya,;A structure theory of Freudenthal-Kantor 
triple systems III,
Mem.Fac.
Sci.Shimane Univ.
23(1989)33-51.
\bibitem{6} 
N.Kamiya,;The construction of all Simple 
Lie algebras over $C$ from 
balanced Freudenthal-Kantor triple systems, 
Contributions to General Algebra
7,Verlag Hoder-Pichler-Tempsky,
Wien,Veralg.G.Teubner,Stutugart(1991)
205-213.
\bibitem{7}
N.Kamiya,;On Freudenthal-Kantor triple 
systems and generalized 
structurable algebras,
Proceeding in International conference of 
nonassociative
algebras and its applications,
Mathematics and Its Applications 303(1994)
198-203,Kluwer Academic Publisher,
\bibitem{8}
N.Kamiya,;On the Peirce decompostions for 
Freudenthal-Kantor triple 
systerns,
Comm.
in 
Alg.
25,(6),(1997),1833-1844.
\bibitem{9}
N.Kamiya, Examples of Peirce decomposition of generalized Jordan triple system of second order,-balanced cases-,
Contemmporary Math., vol.391,AMS.(2005),
Noncommutative Geometry and Representation theory an Mathematical Physics,(ed)Fuchs,
p157-165.
\bibitem{10}
N.Kamiya and S.Okubo, On $\delta$-Lie supertriple systems associated with $(\varepsilon, \delta )$-Freudenthal-Kantor
triple systems, Proc. Edinb. Math.Soc.
43(2000),243-260.
\bibitem{11}
I.L.Kantor and N.Kamiya,
A Peirce decomposition 
for generalized Jordan triple systems of second order,
Comm in Alg.,31(2003),5875-5913.
\bibitem{12} 
I.L.Kantor,;Models of exceptional Lie algebras,
Soviert Math.Dokl.14(1973)
254-258.
\bibitem{13}
I.L.Kantor,;Some generalization of Jordan 
algebras,Trudy Sem.Vektor,
Tensor Ana1.16(1972)407-499(Russian).
\bibitem{14} 
W.Kaup,;Algebraic characterization of symmetric 
complex Banach 
manifoleds,
Math.
Ann.228.
(1977)39-64.
\bibitem{15} 
I.Loos,;Bounded symmetric domains 
and Jordan pairs,Mathematical 
Lectures.lnvine:
University 
of 
California 
at 
Irwine,1977.
\bibitem{16}
K.Meyberg,;Lecture on algebras and triple 
systems,Lecture Notes,the Univ.
ofVirginia,1972.
\bibitem{17}
E.Neher,;On the radical of J*-triples,
Math.Z.211.
(1992)323-332.


\bibitem{1}
B.N.Allison,; A class of nonassociative algebras 
with involution containing class of Jordan algebras, 
Math.Ann., 1978,237(2),133-156.
\bibitem{2} 
N.Jacobson,;Structure and representations of 
Jordan algebras. Amer. Math.
Soc.,Providence.R.1.,1968.
\bibitem{18}
S.Okubo,;lntroduction to Octonion and Other Non-associative
Algebras in
Physics,
Cambridge Univ.Press.1995.
\bibitem{19} 
1.Satake,;Algebraic structure of symmetric 
domains,Princeton University
press.1980.
\bibitem{20}
R.D.Shafer,;An introduction to Non-associative Algebras,Academicpress,
1966.
\bibitem{21}
N.Kamiya and D.Mondoc;
A new class of nonassociative algebras with involution,
Proc.Japan Acad.,84 Ser.A(2008),no5,68-72,
\bibitem{22}
N.Kamiya and S.Okubo,;
A construction of simple Lie superalgebras of certain type from
triple systems,
Bull.Australia Math.Soc. vol.69,(2004) 113-123.
\bibitem{23}N.Kamiya and S.Okubo,;
Construction of Lie superalgebrasD(2,1,$\alpha$),
$G(3)$ and $F(4)$ from
some triple systems.
Proc. Edinburgh Math. Soc. (2003) 46, 87-98.





 	



%\bibitem{}, ``," {\bf }, arXiv:

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath,amsthm,amsfonts}
\usepackage{cite}
\theoremstyle{definition}

\theoremstyle{definition}

\theoremstyle{definition}
 
\begin{document}
\title{\LARGE 
A class of hermitian generalized Jordan \\
triple systems and Chern-Simons gauge theory\\}
\date{}
\author{\Large Noriaki Kamiya \footnote
           {
e-mail address : kamiya@u-aizu.ac.jp}\\
\it Department of Mathematics, University of Aizu\\
\it Aizuwakamatsu, 965-8580, Japan\\ 
\Large Matsuo Sato \footnote
           {
e-mail address : msato@cc.hirosaki-u.ac.jp}\\
\it Department of Natural Science, Faculty of Education, Hirosaki University\\ 
\it Bunkyo-cho 1, Hirosaki, Aomori 036-8560, Japan\\}

\maketitle
\thispagestyle{empty}

\begin{abstract}
We find a class of hermitian generalized Jordan triple systems (HGJTSs) and hermitian $(\epsilon, \delta)$-Freudenthal-Kantor triple systems (HFKTSs). We apply one of the most simple HGJTSs which we find to a field theory, and obtain a typical $u(N)$ Chern-Simons gauge theory with a fundamental matter.
\vskip 3mm
\noindent
AMS classification: 17C50; 17A40; 17B60
\noindent
Keywords: generalized Jordan triple systems, Lie algebras, hermitian operators.
\vskip 0mm
\noindent
APS classification: 11.15.Yc; 11.25.Yb; 11.25.-w 
\noindent
Keywords: 3-algebras, Chern-Simons gauge theories, M-theory
\end{abstract}
\vskip 3mm
{\bf Introduction }
\vskip 2mm
\par
HGJTSs and HFKTSs \cite{KamiyaSato1, KamiyaSato2} are generalizations\footnote{
A different generalization, the so-called differential crossed module, was extensively investigated in \cite{modules1, modules2, modules3}. } of hermitian 3-algebras \cite{Nambu, Filippov, BergshoeffSezginTownsend, deWHN, Yoneya, Minic, BLG1, BLG2, BLG3, BLG4, BLG5, Nogo1, Nogo2, Nogo3, sp1, Iso, sp2, sp3, Lorentz0, Nogo4, Nogo5, Nogo6, Nogo7, text, Nogo8, Nogo9, Nogo10, NishinoRajpoot, GustavssonRey, ABJ, HanadaMannelliMatsuo, IshikiShimasakiTsuchiya, KawaiShimasakiTsuchiya, IshikiShimasakiTsuchiya2, PangWang, DeBellisSaemannSzabo, Jabbari, GomisSalimPasserini, HosomichiLeeLee, Smolin1, Smolin2, Azuma, 3-graded, 5-graded, Palmkvist}, which have played crucial roles in M-theory. The field theories applied with hermitian 3-algebras are the Chern-Simons gauge theories with bi-fundamental matters that describe the effective actions of membranes in M-theory\footnote{The structures of generalized Jordan triple systems were found in the effective actions of the membranes in \cite{Nilsson, Chowdhury}}.  Moreover, 3-algebra models of M-theory itself have been proposed and were studied in \cite{bosonicM, MModel, LorentzianM, Zariski, book, 3BFSS, 3IIB, 4IIB}. Therefore, it is interesting to apply  HGJTSs and HFKTSs to field theories.  
\par
In \cite{KamiyaSato1, KamiyaSato2}, we searched HGJTSs and HFKTSs by adding some terms to the Hermitian 3-algebras. In this paper, we study a class of  HGJTSs and HFKTSs systematically. We apply one of the most simple HGJTSs which we find to a field theory, and obtain a typical $u(N)$ Chern-Simons gauge theory with a fundamental matter. 
\par
Here, we are concerned with algebras and
triple systems that are finite or infinite over a 
complex number field, unless otherwise specified. This paper is an announcement of new results, that is first step of our works for hermitian triple systems and applications to field theories. hence we will give a lot of references \cite{3,4,5,6,7,8,9,10,11,12,13,14,15,16,17}. 
We refer to \cite{1, 2, 18, 19, 20} for nonassociative algebras, and to \cite{21, 22, 23} for Lie superalgebras, for example.
\par
\section{Definitions}
\par
In this section, we summarize definitions of HGJTS and HFKTS given in \cite{KamiyaSato1, KamiyaSato2}.

 A triple system $U$ is a vector space over a field $F$ of characteristic $\neq 2, 3$ with a trilinear map $U \otimes U \otimes U \to U$. In this paper, we are concerned with triple systems over the complex number $\bold{C}$, and denote the trilinear product by $<xyz> \in U$ for $x,y,z \in U$, assuming that $<xyz>$ is $\bold{C}$-linear on $x$ and $z$, and $\bold{C}$-antilinear on $y$. Also we use the notations of two operations $L(x,y) \in$ End $U$ and $K(x,y)\in$ End $U$ with  respect to the triple systems ([1], [2]), where $L(x,y)z=<xyz>$ and $K(x,y)z=<xzy>- \delta <yzx>\ \  (\delta =\pm 1)$.
\vskip 3mm
{\bf Definition.}
\quad
A triple system
$U$ is said to be a *-$(\epsilon, \delta)$-Freudenthal-Kantor triple
system (FKTS) if relations (0)--(iv) are satisfied.
\par
0)\quad  $U$ is a Banach space,
\par
$
i) \ [L(x,y),L(z,w)]=L(<xyz>,w)+\varepsilon L(z,<yxw>) \quad\mbox{(fundamental identity)},
$
\par
\noindent
where $[A,B]=AB-BA$, $A, B \in \mbox{End}\; U, and \ \ \varepsilon =\pm 1$,
\par
$
ii) \ K(<xyz>,w)+K(z,<xyw>)+\delta K(x,K(z,w)y)=0 \quad\mbox{(second order identity)},
$
\par
iii)\quad $<xyz>$ is a ${\bf C}$-linear operator on
$x$ and $z$, and a ${\bf C}$-antilinear operator on $y$,
\par
iv)\quad $<xyz>$ is continuous with respect to a norm,
$||\ ||$ 
that is,
there exists 
\par
\noindent
$K>0$ such that
$$
||<xxx>||\leq K || x ||^{3}
\ {\rm for\ all} \; x\in U.$$
$U$ is called *-generalized Jordan triple system (GJTS) if 0), i), iii), and iv) are satisfied when $\epsilon=-1$ and $\delta=-1$. The condition ii) is called second order identity and *-generalized Jordan triple system of second order is (-1, -1) FKTS. Note that there are many simple Lie algebras (the case of $\delta=1$) and simple Lie superalgebras (the case of $\delta=-1$) constructed from $(\epsilon, \delta)$-Freudenthal-Kantor triple systems ([4,7,8]) because of the second order identity.
\par
Furthermore, a *-$(\epsilon, \delta)$-FKTS  and *-GJTS are said to
be hermitian if
they satisfy the following condition.
\par
v)\quad 
All operators $L(x,y)$
are operators with
a hermitian metric
$$
(x,y)=tr \ L(x,y),$$
that
is,
$(L(x,y)z,w)=(z,L(y,x)w)$
and
$(x,y)=\overline{(y,x)}$.
\par
\vskip 3mm

\section{A class of HGJTSs and HFKTSs}
In this section, we obtain a class of HGJTSs and HFKTSs by making an ansatz $<xyz>=\alpha x \tilde{y} z + \beta z \tilde{y} x + \gamma \bar{y} \tilde{\bar{x}} z$ and explicitly evaluating the conditions i) and ii) in the last section. The hermitian 3-algebra $x \tilde{y} z-z \tilde{y}x$ is in a special case when $\alpha=1$, $\beta=-1$, and $\gamma=0$. By straightforward, but extensive calculations (to be omitted the proof here), we obtain theorems as follows.

\par
\vskip 3mm
{\bf Theorem A.}
\quad
Let $U$ be a hermitian triple system over complex number $\bold{C}$ satisfying the triple product
$<xyz>=\alpha x \tilde{y} z + \beta z \tilde{y} x + \gamma \bar{y} \tilde{\bar{x}} z$ for $\alpha, \beta, \gamma \in \bold{R}$ (real number), $x, y, z \in U$, induced from an associative algebra with an involution $\widetilde{xy} = \tilde{y} \tilde{x}$ and $\tilde{\tilde{x}}=x.$ 

Then we have a necessary and sufficient condition for $U$ to be a hermitian generalized Jordan triple system is that the following hold;

$\gamma( \alpha + \gamma)=0.$

Furthermore, we have for $U$ to have the second order condition ii) is the following holds;

$\alpha=\beta$ if $\delta=1$,

 $\alpha=-\beta$ if $\delta=-1$.

\par
\vskip 3mm

Summarizing, we obtain a class of HGJTSs,
\par
\vskip 3mm
{\bf Proposition B.}
\quad
Let $U$ be a set of $M(n, n; \bold{C})$ with the triple product defined by 

1)$<xyz>$ is a linear combination of $x \tilde{y} z$ and $z \tilde{y}x$,

2) $<xyz>$ is a linear combination of $x \tilde{y} z-\bar{y} \tilde{\bar{x}} z$ and $z \tilde{y}x$,

where $\tilde{x}=\bar{x}^T$, $\bar{}$ is the conjugation of $\bold{C}$, and $x^T$ is the transpose of $x$. Then $U$ are HGJTSs. The hermitian 3-algebra $<xyz>= x \tilde{y} z-z \tilde{y}x$ is in a special case of 1). One of the most simple cases, $<xyz>=x \tilde{y} z$ will be applied to a field theory in the next section.

\par
\vskip 3mm

We also obtain a class of HFKTSs,
\par
\vskip 3mm
{\bf Proposition C.}
\quad
Let $U$ be a set of $M(n, n; \bold{C})$ with the triple product defined by 

a) $<xyz>=x \tilde{y} z + z \tilde{y} x$,

b) $<xyz>=x \tilde{y} z + z \tilde{y} x -\bar{y} \tilde{\bar{x}} z$,

c) $<xyz>=x \tilde{y} z - z \tilde{y} x$,

d) $<xyz>=x \tilde{y} z - z \tilde{y} x +\bar{y} \tilde{\bar{x}} z$,

where $\tilde{x}=\bar{x}^T$, $\bar{}$ is the conjugation of $\bold{C}$, and $x^T$ is the transpose of $x$. Then we have the case of a) and b) are (-1,1)-HFKTSs and the case of c) and d) are (-1,-1)-HFKTSs.

\par
\vskip 3mm
{\bf Remark. }
\quad
For the case of nonassociative and nonhermitian algebras, we note that there is a structureable algebra due to Allison \cite{1} satisfying the triple product
$<xyz>=(x \tilde{y})z+(z\tilde{y})x-(z\tilde{x})y$ and $\widetilde{xy}=\tilde{y}\tilde{x}$, and $\tilde{\tilde{x}}=x$, which is a (-1,1)-Freudenthal-Kantor triple system also this algebra contains a class of commutative Jordan algebras and associative algebras.

We call a hermitian ($\alpha$, $\beta$, $\gamma$) structureable algebra about the hermitian triple system induced from a nonassociative algebra satisfying 
$<xyz>=\alpha (x \tilde{y})z +\beta (z\tilde{y})x+ \gamma(z\tilde{\bar{x}})\bar{y}$. However, the details will be discussed in other paper.

\section{Chern-Simons gauge theory with a fundamental matter}
\par
In this section, we apply one of the most simple HGJTSs to a field theory. We define a 3-bracket, $[x, \overline{y}, z]:=<x y z>$.

We start with
\begin{eqnarray}
S= \int d^3x \mbox{tr}(\! \! \! \!\! \! \! \!\! &&-\bold{D}_{\mu} Z \overline{\bold{D}^{\mu} Z}^T + V(Z) \nonumber \\
 &&+\epsilon^{\mu\nu\lambda}(-\frac{1}{2}A_{\mu \bar{b}c}\partial_{\nu} A_{\lambda \bar{d}a} \overline{T}^{T\bar{d}}[T^c, \overline{T}^{\bar{b}}, T^a ] \nonumber \\
&& \qquad \quad +\frac{1}{3} A_{\mu \bar{d} a} A_{\nu \bar{b} c} A_{\lambda \bar{f} e}[T^c, \overline{T}^{\bar{b}}, T^a] \overline{[T^f, \overline{T}^{\bar{e}}, T^d]})), \label{masteraction}
\end{eqnarray}
where
\begin{equation}
\bold{D}_{\mu} Z=\partial_{\mu} Z -A_{\mu \bar{b} a} [T^a, \overline{T}^{\bar{b}}, Z].
\end{equation}
$Z$ and $A_{\mu}$ are matter and gauge fields, respectively. $V(Z)$ is a potential term. This action is invariant under the transformations generated by the operator $L(x,y)-L(y,x)$. The action (\ref{masteraction}) equipped with  a Lorentzian Lie 3-algebra and a hermitian 3-algebra describes the bosonic parts of the effective actions of supermembranes in M-theory. We apply a HGJTS, 
$[x, \overline{y}, z]=<x y z>=x{\bar y}^{T}z$ in the previous section 
to this action and see what kind of field theory we obtain.

The covariant derivative is explicitly written down as

\begin{equation}
\bold{D}_{\mu} Z= \partial_{\mu} Z -iA_{\mu} Z,\end{equation}
where $A_{\mu}:= -i A_{\mu \bar{b}a} T^a \overline{T}^{T \bar{b}}$ are hermitian matrices, which generate the $u(N)$ Lie algebra. In order to see how gauge transformations act on these fields, let us see how $L(x,y)-L(y,x)$ acts on the coupling between $A_{\mu}$ and $Z$:
\begin{eqnarray}
&&\delta (A_{\mu \bar{b} a}[T^a, \overline{T}^{\bar{b}}, Z]) \nonumber \\
&=& \Lambda_{\bar{d} c} A_{\mu \bar{b} a }([[T^c, \overline{T}^{\bar{d}}, T^a], \overline{T}^{\bar{b}} Z]
-[T^a, \overline{[T^d, \overline{T}^{\bar{c}}, T^b]}, Z]
+[T^a, \overline{T}^{\bar{b}}, [T^c, \overline{T}^{\bar{d}}, Z]] \nonumber \\
&=& 
[i \Lambda, i A_{\mu}]Z
+i A_{\mu}i \Lambda Z \nonumber \\
&=&  i \Lambda i A_{\mu}Z,
\end{eqnarray}
where gauge parameters $\Lambda$ is defined in the same way as  $A_{\mu}$. From this, we obtain gauge transformations,
\begin{eqnarray}
\delta A_{\mu} &=&[i \Lambda, A_{\mu}] \nonumber \\
\delta Z &=& i \Lambda Z.
\end{eqnarray}
Therefore, $A_{\mu}$ transforms as adjoint representations of $u(N)$, whereas $Z$ transforms as a fundamental representation of $u(N)$. As a result, the action can be rewritten in a covariant form and we obtain a typical $u(N)$ Chern-Simons gauge theory with a fundamental matter,
\begin{eqnarray}
S&=&\int d^3x \mbox{tr} (-(\partial_{\mu} Z -iA_{\mu} Z) \overline{(\partial_{\mu} Z -iA_{\mu} Z)}^T +V(Z)\nonumber \\
&& + \epsilon^{\mu\nu\lambda}(\frac{1}{2}A_{\mu} \partial_{\nu} A_{\lambda}
+\frac{i}{3}A_{\mu} A_{\nu} A_{\lambda}). \label{CSaction}
\end{eqnarray}

\vspace*{1cm}

\section*{Acknowledgements}

The work of M.S. is supported in part by Grant-in-Aid for Young Scientists (B) No. 25800122 from JSPS.

\end{document}